OpenStudy (anonymous):

Find the 5th term of the sequence in which t1 = 8 and tn = -3tn-1.

OpenStudy (anonymous):

sorry bro :p

OpenStudy (jdoe0001):

\(\large \bf t_{\color{red}{ 1}}=8\qquad t_{\color{red}{ n}}=-3t_{\color{red}{ n}}-1\qquad t_{\color{red}{ 5}}=-3t_{\color{red}{ 5}}-1\)

OpenStudy (jdoe0001):

hmm

OpenStudy (jdoe0001):

can you post a quick screenshot of the material?

OpenStudy (anonymous):

what kind of material this is the formula if thats what you are asking

OpenStudy (jdoe0001):

the actual question

OpenStudy (jdoe0001):

just a quick screenshot of it

OpenStudy (jdoe0001):

\(\bf t_1=8\qquad t_n={\color{red}{ -3}}t_{n-1}\qquad\textit{comomon ratio}={\color{red}{ -3}}\\ \quad \\ 5^{th}\ term=4^{th}\ term\cdot {\color{red}{ -3}}\\ \quad \\ \begin{array}{llll} t&\\ 1&8\\ 2&8{\color{red}{ (-3)}}\\ 3&8{\color{red}{ (-3)}}{\color{red}{ (-3)}}\\ 4&8{\color{red}{ (-3)}}{\color{red}{ (-3)}}{\color{red}{ (-3)}}\\ 5&8{\color{red}{ (-3)}}{\color{red}{ (-3)}}{\color{red}{ (-3)}}{\color{red}{ (-3)}}\implies 8{\color{red}{ (-3)}}^{n-1} \end{array}\\ \quad \\ t_{\color{red}{ 5}}=t_1(-3)^{{\color{red}{ 5}}-1}\)

OpenStudy (anonymous):

t5=t1(−3)5−1 this is the answer?